value 0.50 points A company can buy a machine that is expected to have a three-year life and a $39,000 salvage value. The machine will cost $1,836,000 and is expected to produce a $209,000 after-tax net Income to be recelved at the end of each year. If a table of present values of 1 at 12% shows values of 0.8929 for one year, 0.7972 for two years, and 0.7118 for three years, what is the net present value of the cash flows from the Investment discounted at 12%? $132,496 $602,895 O $644,138 O $721,463 O $1,968,496
Answer #1

Cash inflow = 209000+(1836000-39000/3) = 808000

Net present value = Present value of cash inflow-Present value of cash outflow

= (808000*2.4019+39000*.7118)-1836000

Net present value = 132496

so answer is a) 132496